13 August 2014: Wednesday

LDAP:

  • Issues!
    • Two former students: dancing done.
    • A singleton match with a different last name.  Off to the provider to get them to check it out.
  • More work with the new TWS/Maestro java stuff that really doesn’t like our LDAP server.

BACKUPS:

  • Tape clones completed yesterday.  Disk clones have 41TB left to go (egad..).  CoM/IS clones are not started yet – their weekend full has not yet completed.
  • Data shredding of the disk array used in the testing of Veeam has completed, the array and the system it was attached to have been powered off.
  • CoM/IS backup finally completed.  CoM/IS clones started at noon.

 

Tags:

Leave a Reply

You must be logged in to post a comment.